OpenStudy (anonymous):

In 1920, the record for a certain race was 45.4 seconds. In 1980 it was 44.2 seconds. Let R(t) = the record in the race and t=the number of years since 1920. R(t) =? What is the predicted record for 2003? 2006? And in what year will the predicted record be 43.64 seconds?

OpenStudy (a_clan):

R(t)= 45.4 - 0.02t for 2003, t= (2003-1920)=83 So, R(83) = 45.4 - 0.02*83 = 43.74 Similarly for 2006, t= 86 last part of the question, 43.64 = 45.4 - 0.02t Add this t to 1920